OpenStudy (anonymous):

Consider the parabola y = 5x + x^2 a) Find the slope of the tangent line to the parabola at the point (1, 4). b) Find the equation of the tangent line of part (a)

OpenStudy (anonymous):

so the parabola is \[y = 5x + x ^{2}\]

OpenStudy (loser66):

Where are you stuck?

OpenStudy (anonymous):

the slope of the tangent line would i use the slop formula?

OpenStudy (loser66):

yes

OpenStudy (anonymous):

so \[y- y ^{1} = m(x-x ^{1^{}})\]

OpenStudy (anonymous):

and i would input the tangent line for y1 and x1

OpenStudy (gorv):

just differentiate y with respect to x

OpenStudy (gorv):

dy/dx=5+2x at (1,4) dy/dx=5+2*1=7

OpenStudy (gorv):

now you have slope and ponit it pass through

OpenStudy (gorv):

use y=mx+c put point in equation 4=7*1+c c=-3 so tangent y=7x-3

OpenStudy (aum):

The formula you used is for finding slope of a straight line. Here it is curve. So you will have to find the derivative and then plug in x value to find the slope.

OpenStudy (loser66):

@holls622 When you said "the slope of the tangent line would i use the slop formula?", I assumed you know how to find the "slop formula" meaning